INAF queue observing, 2016 Jan 30-31

Observer: F. Cusano, A. Rossi
Telescope Operator: S. Allanson
Support Astronomer: Olga Kuhn
Instrument: LBC

Summary

At the sunset clouds. The dome stay closed for high wind gusts (25 m/s)

Night duration (12\xB0 el twilight) 19:20-6:10 ( 02:20-13:10 UT) = 10 h 50m

Observing time: 5h 30m (55%)

Weather time loss: 5 h (45%)

technical time loss * 0h* 0m (%)

Data Summary

Project_RUNSorted ascending Objects Filter ExpTime Notes
Antoniucci NY_Ori dual grating 15m with imaging
Antoniucci V512 Per dual grating 30m with imaging
Antoniucci V1118 Ori dual grating 20m high wind
Castellano MACSJ0717   15m OB interrupted seeing 2 out of constraints
Guaita M1149 Br_gamma 0.5  
MODSBias       bias8k, bias3k
MODSCalib       flats 5" red ok, lamps for 1" dual ok, flat 0.8" dual ok, flat 1" NOT OK!

Details

00:37 Start with MODS calibrations
  • slit flat for standard. OB slitflat_5as_r mods1r.20160131.0001-0003
  • slit arclamp for 1", dual grating mods1b.20160131.0001-0009, mods1r.20160131.0004-0012
  • slit flat for 0.8" and 1", dual grating mods1b.20160131.0010-0018, mods1r.20160131.0013-0021, the last files maybe are not good because the TO opened the dome to start observations

01:35 the TO opens the dome
  • we start with LBCs

01:42 we slew to copointing of Dotto 285179
  • the target is in the direction of the wind, we change target

01:44 we slew to copointing of Castellano, Castellano_MACSJ0717_pointing.ob
  • 1:48 dofpia, problem with the collimation we are giving correction by hand
  • 2:22 reached convergence, we run lbcrangebal

2:28 science of Castellano, executing Castellano_MACSJ0717_UV_7.ob
  • seeing 2.0" on the images, out of constrain we change program

2:44 reconfiguring for MODS

03:31 slew to target NY_Ori for Antoniucci
  • seeing 2" on the guider at 3:31
  • acq mods1r.20160131.0022-0024
  • spec mods1b.20160131.0020-0023, mods1r.20160131.0025-0028
  • seeing 1.7" at 3:42
  • seeing 1.9" at 3:53

04:02 slew to V512 Per
  • acq mods1r.20160131.0029-0031
  • spec mods1b.20160131.0024-0027, mods1r.20160131.0032-0035
  • seeing 1.5" at 4:12
  • seeing variable between 1.5 and 2.2"
  • imaging mods1b.20160131.0028-0030, mods1r.20160131.0036-0038

05:02 slewing to V1118 Ori
  • seeing 1.9" at 5:13, wind at 20m/s
  • acq mods1r.20160131.0039-0041
  • spec mods1b.20160131.0031-0034, mods1r.20160131.0042-0045
  • seeing 1.6" at 5:21
  • 5:40 lost guide star due to high wind we can not performe imaging, spectra are ok

05:43 switch to LUCI
  • average dimm seeing 2"
  • problem with the rotator of the telescope, the TO is trying to solve the problem with technicians

06:10 DOME Closed due to wind>20m/s

MODS BIAS
  • script bias8k mods1b.20160131.0035-0039, mods1r.20160131.0046-0050
  • script bias3k mods1b.20160131.0039-0044, mods1r.20160131.0051-0055

09:00 Dome open wind below 20 m/s

09:01 Slew to Guaita CD, target M1149
  • pointing, start OB at 09:10, seeing 1.2" on guider
  • scheduler problem, luci2.20160131.0001-002 NOT GOOD: mask is in FPU.
  • 09:17 start again ima luci2.20160131.0003-0032
  • seeing 1.2" on image at 09:25, keep going
  • seeing 1" on image at 9:31
  • ima luci2.20160131.0019 not good, passing cirrus
  • seeing 1,1" on image at 9:48
  • variable seeing, check image
  • seeing 1.3" on image at 09:58
  • after image luci2.20160131.0028 seeing too variable, high wind

10:05 Slew to Standard for De Barros

10:06 DOME Closed due to wind>20m/s

10:15 running MODS calibration
  • flat for 0.8" and 1" slit, to be sure, because some of the ones taken in the afternoon were with the dome opened

12:00 Wind still high, no chances to open, we finish here

 
Topic revision: r1 - 31 Jul 2024, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback